OpenStudy (anonymous):

How do we factorize this? I'm confused. sin^2 (x) + 2cos (x) + 2.

OpenStudy (aaronq):

are you trying to solve for x? you probably gotta use some identities

OpenStudy (tkhunny):

And you may wish to remember that \(\sin^{2}(x)+\cos^{2}(x) = 1\)

OpenStudy (anonymous):

which is another way of saying you can replace \(\sin^2(x)\) by \(1-\cos^2(x)\) and end up only with cosine

OpenStudy (anonymous):

you got that?

OpenStudy (anonymous):

Yes sir. Thank you all for clarification. Such simple answer, yet I was hung up.

OpenStudy (anonymous):

yw